(Can someone help me with this question please?, thanks)
Solve the system of equations by graphing. Then classify the system.
3x-9y=36
2x-6y=-12
What is the solution of the system?
a)The solution is___?
b)There are infinitely many solutions
c)There is no solution
Is the system consistent or inconsistent?
a)inconsistent
b)consistent
Are the equations dependent or independent?
a)dependent
b)independent

Answer by ewatrrr(24785)

 
Hi,
3x-9y=36 |Multiply by 2 to eliminate x
2x-6y=-12 |Multiply by -3 to eliminate x
6x-18y=72
-6x+18y=36
0 ≠ 108 There is NO soluion
A consistent system has at least one solution.
An inconsistent system has no solutions.****
An independent system has exactly one solution.
A dependent system has infinitely many solutions.